OpenStudy (thomas5267):

Please check my steps:

OpenStudy (thomas5267):

\[ \begin{align*} \int_0^\infty tf(t)dt &= \lambda\int_0^\infty te^{-\lambda t}dt\\ &=\lambda\left[\left(-\frac{1}{\lambda}te^{-\lambda t}\right)_0^\infty-\int_{0}^{\infty}-\frac{1}{\lambda} e^{-\lambda t}dt\right]\\ &=\lambda\left[-\frac{1}{\lambda}\left(te^{-\lambda t}\right)_0^\infty + \frac{1}{\lambda}\int_{0}^{\infty}e^{-\lambda t}dt \right]\\ &=\lambda\left[-\frac{1}{\lambda}\left(\lim_{t\to\infty}\left(\frac{t}{e^{\lambda t}}\right)-0\right)-\frac{1}{\lambda^2}\left(e^{-\lambda t}\right)_0^\infty\right]\\ &=-\lim_{t\to\infty}\left(\frac{t}{e^{\lambda t}}\right)-\frac{1}{\lambda}\left(\lim_{t\to\infty}\left(e^{-\lambda t}\right)-1\right)\\ &=-\lim_{t\to\infty}\left(-\frac{1}{\lambda e^{-\lambda t}}\right)+\frac{1}{\lambda}\\ &=\frac{1}{\lambda} \end{align*} \]

hartnn (hartnn):

in 2nd last step, that \(\Large e^{-\lambda t }\) should be in numerator. rest all steps are correct :)

OpenStudy (thomas5267):

Ah! The second last step is wrong! Is this correct though? \[ \begin{align*} -\lim_{t\to\infty}\left(\frac{t}{e^{\lambda t}}\right)-\frac{1}{\lambda}\left(\lim_{t\to\infty}\left(e^{-\lambda t}\right)-1\right)&=-\lim_{t\to\infty}\frac{1}{\lambda e^{\lambda t}}+\frac{1}{\lambda}\\ &=\frac{1}{\lambda} \end{align*} \]

hartnn (hartnn):

oh, you used L'Hopital's rule, right ?? then there is only sign error in 2nd last step, should be \(λe^{λt}\) in the denominator

hartnn (hartnn):

yep, everything elase is correct :)

OpenStudy (thomas5267):

One last check please: \[ \begin{align*} \int_0^\infty tf(t)dt &= \lambda\int_0^\infty te^{-\lambda t}dt\\ &=\lambda\left[\left(-\frac{1}{\lambda}te^{-\lambda t}\right)_0^\infty-\int_{0}^{\infty}-\frac{1}{\lambda} e^{-\lambda t}dt\right]\\ &=\lambda\left[-\frac{1}{\lambda}\left(te^{-\lambda t}\right)_0^\infty + \frac{1}{\lambda}\int_{0}^{\infty}e^{-\lambda t}dt \right]\\ &=\lambda\left[-\frac{1}{\lambda}\left(\lim_{t\to\infty}\left(\frac{t}{e^{\lambda t}}\right)-0\right)-\frac{1}{\lambda^2}\left(e^{-\lambda t}\right)_0^\infty\right]\\ &=-\lim_{t\to\infty}\left(\frac{t}{e^{\lambda t}}\right)-\frac{1}{\lambda}\left(\lim_{t\to\infty}\left(e^{-\lambda t}\right)-1\right)\\ &=-\lim_{t\to\infty}\left(\frac{1}{\lambda e^{\lambda t}}\right)+\frac{1}{\lambda} &\text{L'Hôpital's rule}\\ &=\frac{1}{\lambda} \end{align*} \]

hartnn (hartnn):

yep. all are correct, good work! :)